For the next fiscal​ year, you forecast net income of or the next fiscal​ year, you forecast net income of $ 49400 and ending assets of $ 505600 Your​ firm's payout ratio is 9.6 %.Your beginning​ stockholders' equity is $ 299 comma 900and your beginning total liabilities are $119900Your​ non-debt liabilities such as accounts payable are forecasted to increase by $10300 Assume your beginning debt is $ 101900 What amount of equity and what amount of debt would you need to issue to cover the net new financing in order to keep your​ debt-equity ratio​ constant?
Answer #1

Answer:-

Step-1

Formula:

Beginning debt equity ratio constant = beginning stock holder equity/ beginning debt

Beginning debt equity ratio constant= 299900/(101900+119900)

= 299,900/221800=1.35

Step-2

formula:

Final debt equity ratio constant = Fina stock holder equity/final debt

Final debt equity ratio constant = Fina stock holder equity/final debt=1.35

Retained earnings = 49400-(49400*0.096)=44657.6

Final stock holder equity = 271267.8+44657.6=315925.4

Final debt equity ratio constant= 315925.4/(505600-271267.8)=1.35